Add the following unsigned binary numbers as shown.

a) 01000100 + 10111011 =

b) 01011011 + 00011111 =

c) 10101100 + 00100100 =

showing work would help with future problems
Answer #1

a) 01000100 + 10111011 =

01000100
+ 10111011
---------------
11111111

here we have to add the first number with second number from left most to rightmost direction. In this question there is no overflow
in right most digit

b) 01011011 + 00011111 =   
11111
01011011
+ 00011111
---------------
01111010

here we have added carry by passing to next digit.

c) 10101100 + 00100100 =
1 11
10101100
+ 00100100
----------------
11010000


here we have added carry by passing to next digit.